An alumina refinery is using process carbon dioxide (CO2) from a nearby fertilizer plant to reduce the alkalinity of its bauxite residue, thus reducing environmental risks and significant ongoing management related to bauxite residue storage areas while also leaving options open for additional processing of the residue into other useful products at a future stage. The alumina refinery uses 10 kg CO2 for reducing alkalinity of 1 tonne bauxite residue, while the refinery produces 50,000 tonnes of alumina per year. The bauxite which is refined to alumina is low grade bauxite (40%). The refinery is expected to reduce carbon taxes for sequestering CO2 from the neighbouring plant. The Government has just implemented Carbon tax of $20/tonne of CO2 emitted. Calculate the economic and environmental benefits from this industrial symbiosis. (4 marks)
Treatment 10 kg CO2 /tonne of Bauxite residue
Production 50,000 tonnes/year
Bauxite in alumina 40%
Needs 50,000/0.4 or 125,000 tonnes Bauxite to produce 50,000 tonne of alumina
Therefore bauxite residue is (125,000-50,000) tonnes or 75,000 tonnes
Environmental benefit = 10 * 75,000 = 750,000 kg = 750 tonnes of CO2-e
C- tax $20/tonne of bauxite residue
Economic benefit is 750 * 20 = $15,000
Answer #1

Annual Alumina production = 50000 tonnes

Bauxite (Alumina/Residue) % = 40%=0.4

i.e from 100 tonnes of Bauxite, only 40 Tonnes are alumina and remaining 60 tonnes are residue.

Therefore,Annual demand for Bauxite = 50000 ×(100/40) = 1,25,000 tonnes .

Of this 1,25,000 tonnes of Bauxite, only 50,000 tonnes are Alumina and remaining 75,000 tonnes are residue,which needs to be reduced by CO2.

As per given conditions,

10 kg CO2 is required to reduce 1 ton of Residue.

So, for 75,000 tonnes of Residue, Co2 required is 7,50,000 kg or 750 Tonnes of Co2 ( Environmental benefit)

Tax levied is $20 per ton of Co2

Therefore for 750 tonnes of Co2, Money required for taxes = 20×750= $ 15,000 (Economic benefit)